OpenStudy (anonymous):

Which number line shows the given inequality? k > 4.5 A. B. C. D.

OpenStudy (igreen):

If it has >, it goes to the right. If it has <, it goes to the left. If it has < or >, there's an open dot. If it has \(\le\) or \(\ge\), there's a closed dot.

OpenStudy (igreen):

So which one do you think?

OpenStudy (anonymous):

tho's are the answers ^

OpenStudy (anonymous):

it is in A B C and D order

OpenStudy (igreen):

Okay..so which one do you think?

OpenStudy (anonymous):

I was thinking C or D

OpenStudy (anonymous):

but not sure

OpenStudy (anonymous):

I know it is one of the two C or D

OpenStudy (igreen):

D is incorrect because it has a closed dot.

OpenStudy (igreen):

It's \(\sf k > 4.5\) and not \(\sf k \ge 4.5\), right?

OpenStudy (anonymous):

ya that's what get's me confused is the Dot's

OpenStudy (anonymous):

thanks iGreen